OpenStudy (anonymous):

Find the inverse of the function. f(x) = the cube root of quantity x divided by eight. - 4

OpenStudy (anonymous):



OpenStudy (anonymous):

This is my answer f-1(x) = 8(x + 4)^3 , I just need to verify.

OpenStudy (anonymous):

actually, the answer is \[f^-1(x)=\frac{ 2 }{ \sqrt[3]{x} -8 }\]

OpenStudy (anonymous):

sorry it took so long for me to respond, still getting used to the equation writer system

OpenStudy (anonymous):

It's okay, but that is not one of my options.

OpenStudy (anonymous):

f-1(x) = [8(x + 4)]3 f-1(x) = 8(x + 4)3 f-1(x) = 8(x3 + 4) f-1(x) = 24(x + 4)

OpenStudy (anonymous):

anyway, the easy way for you to check your answer is to take the equation and put it under 1, like this: \[\frac{ 1 }{ \sqrt[3]{x/8} -4 }\]

OpenStudy (anonymous):

I understand, it's just that your answer does not compare to the options at all.

OpenStudy (anonymous):

yes, i noticed, but all of those answers are easily proved incorrect simply by giving x a value and working the equation out from there

OpenStudy (anonymous):

i tested this 3 times, i made x = 4 then i tried x = 25 then i tried x=64 all of these show that my answer is correct and the all of the answer to the question are incorrect

OpenStudy (anonymous):

tell me if im looking at this correctly is it? \[\sqrt[3]{x/8}-4\]

OpenStudy (anonymous):



OpenStudy (anonymous):

yes, thats what i thought

OpenStudy (anonymous):

Where you have the a, it should be a 3.

OpenStudy (anonymous):

it is a 3, it should be anyway i wrote it like that